# Fairness-aware Contextual Dynamic Pricing with Strategic Buyers

### v1

@@ 7 Summary and Future Directions
Appendix S.2 Proof of Theorem 1
We assume that the dimension of the features is $d=1$, and the expected demands of group 0 and group are $\mathbb{E}y_{0t}=2+x_{t}-p_{0t}$ and $\mathbb{E}y_{1t}=2+x_{t}-2p_{1t}$, respectively. We assume $x_{t}\sim$ Unif(-1/2, 1/2) and set $q=1/2,\delta=1/4$. By (3), the optimal prices for group 0 and group 1 under the fairness constraint are $p_{0}^{*}(x_{t})=x_{t}/3+5/6$ and $p_{1}^{*}(x_{t})=x_{t}/3+7/12$, respectively. We denote $p_{jt}^{*}=p_{jt}^{*}(\bm{x}_{t})$ and $p_{jt}=p_{jt}(\bm{x}_{t})$ for $j=0,1.$. The revenue of the optimal pricing policy at time $t$ under the fairness constraint is
$\frac{1}{2}[R_{0}(p_{0t}^{*},x_{t})+R_{1}(p_{1t}^{*},x_{t})]$ $=\frac{1}{2}[p_{0t}^{*}(x_{t}-p_{0t}^{*})+p_{1t}^{*}(x_{t}-2p_{1t}^{*})]$ (S2)
@@
$\leq\sum_{t=1}^{T}\mathbb{E}\left[\frac{x_{t}^{2}}{6}+\frac{17x_{t}}{24}+\frac{35}{48}-\frac{(2+x_{t})^{2}}{6}\right]$
$\leq\frac{T}{16}.$
@@
$\leq\mathbb{E}\bigg{[}\frac{\|Z_{t}^{\top}\bm{\epsilon}_{t}\|_{2}^{2}I(\lambda_{min}(Z_{t}^{\top}Z_{t})\geq\lambda_{0}t/2)}{\lambda_{min}^{2}(Z_{t}^{\top}Z_{t})}\bigg{]}+2(a_{max}^{2}+b_{max}^{2})\mathbb{P}\bigg{(}\lambda_{min}^{2}(Z_{t}^{\top}Z_{t})<\frac{\lambda_{0}t}{2}\bigg{)}$
@@
$=\widehat{p}_{0}(\bm{x})-p_{0}(\bm{x})+p_{1}(\bm{x})-\widehat{p}_{1}(\bm{x})+\frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}}-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}}-\delta+\delta$
@@
$\widehat{\delta}-C_{0}\leq\widehat{p}_{0}(\bm{x})-p_{0}(\bm{x})+p_{1}(\bm{x})-\widehat{p}_{1}(\bm{x})-c_{\delta}\sqrt{\frac{\log T_{0}}{T_{0}}}+\delta-C_{0}$ (S9)
@@
$\widehat{\delta}-C_{0}=\widehat{p}_{0}(\bm{x})-p_{0}(\bm{x})+p_{1}(\bm{x})-\widehat{p}_{1}(\bm{x})+\delta-C_{0}$ (S10)
@@
At time $t$, buyers learn the price difference $\widehat{\delta}$ using $t-1$ samples. By Assumption 3, with at least probability $1-2\eta_{t-1}$, we have $|\widehat{p}_{0}(\bm{x})-p_{0}(\bm{x})|\leq\mathcal{E}_{\mathcal{P},\eta_{t-1}}(t-1)$ and $|p_{1}(\bm{x})-\widehat{p}_{1}(\bm{x})|\leq\mathcal{E}_{\mathcal{P},\eta_{t-1}}(t-1)$. Therefore, with at least probability $1-2\eta_{t-1}$,
@@
Assume $10<\eta<$ 1. Let $T_{0}$ be the length of the exploration phase and $q$ be the proportion of buyers in group 0. $\widehat{\bm{\theta}}_{0}$ and $\widehat{\bm{\theta}}_{1}$ are obtained by (2). Under Assumptions 1 and 2, We have with probability at least $1-\eta-(d+2)(e/2)^{-\frac{\lambda_{0}qT_{0}}{2L}}$,
@@
$\|\widehat{\bm{\theta}}_{1k}-\bm{\theta}_{1}\|_{2}^{2}\leq\frac{8(d+2)z_{max}^{2}\sigma_{\epsilon}^{2}\log(2/\eta)}{\lambda_{0}^{2}(1-q)T_{0}},$
@@
Recall that $Z_{t}=(z_{1},\cdots,z_{t})^{\top}\in\mathbb{R}^{t\times(d+2)}$ with $z_{t}=(1\ p_{t}\ \bm{x}_{t}^{\top})^{\top}\in\mathbb{R}^{d+2}$ from the exploration phase. We slightly abuse the notation and let $z_{ij}$ be the $(i,j)$-th elements of $Z_{t}$. Under Assumption 1, we have $|z_{ij}|\leq\mathop{\max}\{1,x_{max},B\}:=z_{max}$. Noting that $\epsilon_{i}$ is bounded independent $\sigma_{\epsilon}^{2}$-sub-Gaussian variable with mean zero, and $z_{ij}$ is independent from $\epsilon_{i}$, we know that $z_{ij}\epsilon_{i}$ is zero-mean bounded random variable with variance at most $z_{max}^{2}\sigma_{\epsilon}^{2}$. By Lemma S10, for $10<\eta_{j}<$ 1, we have
$~{}~{}~{}~{}\mathbb{P}\bigg{(}\bigg{|}\sum_{i=1}^{t}z_{ij}\epsilon_{i}\bigg{|}<z_{max}\sigma_{\epsilon}\sqrt{2t\log(2/\eta_{j})}\bigg{)}$ (S16)
$=1-\mathbb{P}\bigg{(}\bigg{|}\sum_{i=1}^{t}z_{ij}\epsilon_{i}\bigg{|}\geq z_{max}\sigma_{\epsilon}\sqrt{2t\log(2/\eta_{j})}\bigg{)}$
$=1-\mathbb{P}\bigg{(}\sum_{i=1}^{t}z_{ij}\epsilon_{i}\geq z_{max}\sigma_{\epsilon}\sqrt{2t\log(2/\eta_{j})}\bigg{)}-\mathbb{P}\bigg{(}-\sum_{i=1}^{t}z_{ij}\epsilon_{i}\geq z_{max}\sigma_{\epsilon}\sqrt{2t\log(2/\eta_{j})}\bigg{)}$
$\geq 1-\eta_{j}.$
Let $10<\eta<$ 1 and $\eta_{j}=\eta/(d+2)$ for $j=1,\cdots,(d+2)$. By (S16), with probability at least $1-\eta$, we have
$\bm{\epsilon}_{t}^{\top}Z_{t}Z_{t}^{\top}\bm{\epsilon}_{t}$ $=\sum_{j=1}^{d+2}\bigg{(}\sum_{i=1}^{t}z_{ij}\epsilon_{i}\bigg{)}^{2}\leq 2(d+2)z_{max}^{2}\sigma_{\epsilon}^{2}t\log(2/\eta).$ (S17)
@@
$\|\widehat{\bm{\theta}}-\bm{\theta}\|_{2}^{2}\leq\frac{\bm{\epsilon}_{t}^{\top}Z_{t}Z_{t}^{\top}\bm{\epsilon}_{t}}{\lambda^{2}_{min}(Z_{t}^{\top}Z_{t})}\leq\frac{8(d+2)z_{max}^{2}\sigma_{\epsilon}^{2}\log(2/\eta)}{\lambda_{0}^{2}t},$ (S18)
with probability at least $1-\eta-(d+2)(e/2)^{-\frac{\lambda_{0}t}{2L}}$. Since the length of the exploration phase is $T_{0}$, and the proportion of buyers from group 0 is $q$, the numbers of samples used to estimate $\bm{\theta}_{0}$ and $\bm{\theta}_{1}$ are $qT_{0}$ and $(1-q)T_{0}$, respectively. Plugging $t=qT_{0}$ and $t=(1-q)T_{0}$ into (S18), respectively, we can obtain the estimation errors of $\bm{\theta}_{0}$ and $\bm{\theta}_{1}$. ∎
By Lemma S5 with $\eta=2/T_{0}$, we have with probability at most $\frac{4}{T_{0}}+(d+2)[(e/2)^{-\frac{\lambda_{0}qT_{0}}{2L}}+(e/2)^{-\frac{\lambda_{0}(1-q)T_{0}}{2L}}]$,
@@
$J_{0}:=\mathbb{E}(reg_{t})$ $\leq 2B(a_{max}B+b_{max}x_{max})\left\{\frac{4}{T_{0}}+(d+2)[(\frac{e}{2})^{-\frac{\lambda_{0}qT_{0}}{2L}}+(\frac{e}{2})^{-\frac{\lambda_{0}(1-q)T_{0}}{2L}}]\right\}$ (S21)
@@
$p_{1t}=\left\{\begin{aligned} &-\frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}},&\text{if}\ \ \frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}}-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}}\leq\delta-c_{\delta}\sqrt{\frac{\log T_{0}}{T_{0}}},\\ &\widehat{\bm{\gamma}}_{1}^{\top}\widetilde{\bm{x}}_{t}-\delta+\widehat{\gamma}_{2},&\text{if}\ \ \frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}}-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}}\geq\delta+c_{\delta}\sqrt{\frac{\log T_{0}}{T_{0}}}.\end{aligned}\right.$
@@
$p_{0t}^{\prime}=\left\{\begin{aligned} &-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}},\ &\text{if}\ \ \frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}}-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}}\leq\delta-c_{\delta}\sqrt{\frac{\log T_{0}}{T_{0}}},\\ &\widehat{\bm{\gamma}}_{1}^{\top}\widetilde{\bm{x}}_{t}+\widehat{\gamma}_{2},\ &\text{if}\ \ \frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}}-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}}\geq\delta+c_{\delta}\sqrt{\frac{\log T_{0}}{T_{0}}}.\end{aligned}\right.$
The regret of our policy depends on the probability $\mathbb{P}(\widehat{\delta}\leq C_{0})$. We define the historical information up to time $t$ as $\mathcal{H}_{t}=\{\bm{x}_{1},\cdots,\bm{x}_{t},y_{1},\cdots,y_{t},p_{1},\cdots,p_{t}\}$. We also define ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}\cup\{\bm{x}_{t+1}\}$ as the filtration including the feature $\bm{x}_{t+1}$. The expected regret at time $t$ in the exploitation phase is
@@
$=\mathbb{E}\{q\alpha_{0}(p_{0t}^{*2}-p_{0t}^{\prime 2})+(1-q)\alpha_{1}(p_{1t}^{*2}-p_{1t}^{2})+q\bm{\beta}_{0}^{\top}\widetilde{\bm{x}}_{t}(p_{0t}^{*}-p_{0t}^{\prime})+(1-q)\bm{\beta}_{1}^{\top}\bm{x}_{t}(p_{1t}^{*}-p_{1t})\}$
@@
$\leq a_{max}\mathbb{E}(p_{0t}^{*}-p_{0t}^{\prime})^{2}$
The fourth equality is from $p_{1t}^{*}=p_{0t}^{*}-\delta$ and $p_{1t}=p_{0t}^{\prime}-\delta$. The last second equality is due to $2(q\alpha_{0}+(1-q)\alpha_{1})p_{0t}^{*}=2(1-q)\alpha_{1}\delta-(q\bm{\beta}_{0}+(1-q)\bm{\beta}_{1})^{\top}\widetilde{\bm{x}}_{t}$ from (3). We now upper bound the price difference between the optimal policy and our policy. By (4), we rewrite the pricing parameters as
@@
$R_{0}(p^{*}_{0t})-R_{0}(p_{0t}^{\prime})$ $=p^{*}_{0t}(\alpha_{0}p^{*}_{0t}+\bm{\beta}_{0}^{\top}\widetilde{\bm{x}}_{t})-p_{0t}^{\prime}(\alpha_{0}p_{0t}^{\prime}+\bm{\beta}_{0}^{\top}\widetilde{\bm{x}})$
@@
$=T_{0}\bar{B}+\frac{c^{\prime}_{4}T}{T_{0}}+2\sum_{t=T_{0}}^{T}(\frac{c^{\prime}_{4}}{T_{0}}+q\bar{B})\eta_{t-1}$
@@
$=\sqrt{\left[\frac{c_{2}(d+2)+c_{3}}{q(1-q)}\right]T}+(c_{4}+c_{5}q)\sum_{t=2}^{T}\eta_{t-1}$
for some positive constants $c_{2},c_{3},c_{4}$ and $c_{5}$.
Appendix S.6 Proof of Theorem 3
@@
    The third equality is from $p_{0}^{*}-p_{1}^{*}=\delta$ and $p_{0}-p_{1}=\delta$. The fourth equality is due to $\delta=1/4$ and $16\alpha p_{0}^{*}=\frac{7\alpha}{2}-$ 1 derived from $p_{0}^{*}=\frac{7}{12}-\frac{1}{6\alpha}$. the last line is from $\alpha\in[-1/2,-1/5]$. Similarly, we can obtain $R_{0}(p_{0}^{*},\alpha)-R_{0}(p_{0},\alpha)+R_{1}(p_{1}^{*},\alpha)-R_{1}(p_{1},\alpha)\geq\frac{3}{5}(p_{1}^{*}-p_{1})^{2}$.
@@
$\mathbf{X}_{k}\succeq\bm{0}\ \textit{and}\ \lambda_{max}(\mathbf{X}_{k})\leq L\ \textit{almost surely.}$.

### v2

@@ Acknowledgment
S.9.2 Proof of Theorem 1
To begin the proof, we first construct an instance. We assume that the dimension of the features is $d=1$, and the expected demands of group 0 and group are $\mathbb{E}y_{0t}=2+x_{t}-p_{t}$ and $\mathbb{E}y_{1t}=2+x_{t}-2p_{t}$, respectively. We assume $x_{t}\sim$ Unif(-1/2, 1/2) and set $q=1/2,\delta=1/4$. By (3), the optimal prices for group 0 and group 1 under the fairness constraint are $p_{0}^{*}(x_{t})=x_{t}/3+5/6$ and $p_{1}^{*}(x_{t})=x_{t}/3+7/12$, respectively. We denote $p_{jt}^{*}=p_{j}^{*}(\bm{x}_{t})$ and $p_{jt}=p_{j}(\bm{x}_{t})$ for $j=0,1$. The expected revenue of the optimal pricing policy at time $t$ under the fairness constraint is
$\frac{1}{2}[R_{0}(p_{0t}^{*},x_{t})+R_{1}(p_{1t}^{*},x_{t})]$ $=\frac{1}{2}[p_{0t}^{*}(2+x_{t}-p_{0t}^{*})+p_{1t}^{*}(2+x_{t}-2p_{1t}^{*})]$ (S6)
@@
$\geq\sum_{t=1}^{T}\mathbb{E}\left[\frac{x_{t}^{2}}{6}+\frac{17x_{t}}{24}+\frac{35}{48}-\frac{(2+x_{t})^{2}}{6}\right]$
$\geq\frac{T}{16}.$
@@
$\leq\mathbb{E}\bigg[\frac{\|Z_{t}^{\top}\bm{\epsilon}_{t}\|_{2}^{2}I(\lambda_{min}(Z_{t}^{\top}Z_{t})\geq\lambda_{0}t/2)}{\lambda_{min}^{2}(Z_{t}^{\top}Z_{t})}\bigg]+2(a_{max}^{2}+b_{max}^{2})\mathbb{P}\bigg(\lambda_{min}(Z_{t}^{\top}Z_{t})<\frac{\lambda_{0}t}{2}\bigg)$
@@
$=\widehat{p}_{0}(\bm{x})-p_{0}(\bm{x})+p_{1}(\bm{x})-\widehat{p}_{1}(\bm{x})+\frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}}-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}}$
@@
$\widehat{\delta}-C_{0}\leq\widehat{p}_{0}(\bm{x})-p_{0}(\bm{x})+p_{1}(\bm{x})-\widehat{p}_{1}(\bm{x})-c_{\delta}\sqrt{\frac{\log T_{0}}{T_{0}}}+\delta-C_{0}.$ (S13)
@@
$\widehat{\delta}-C_{0}=\widehat{p}_{0}(\bm{x})-p_{0}(\bm{x})+p_{1}(\bm{x})-\widehat{p}_{1}(\bm{x})+\delta-C_{0}.$ (S14)
@@
At time $t$, buyers learn the price difference $\widehat{\delta}$ using t1t-1 samples. By Assumption 3, with at least probability $1-2\eta_{t-1}$, we have $|\widehat{p}_{0}(\bm{x})-p_{0}(\bm{x})|\leq\mathcal{E}_{\mathcal{P},\eta_{t-1}}(t-1)$ and $|p_{1}(\bm{x})-\widehat{p}_{1}(\bm{x})|\leq\mathcal{E}_{\mathcal{P},\eta_{t-1}}(t-1)$. Therefore, with at least probability $1-2\eta_{t-1}$,
@@
Assume $0<\eta<1$. Let $T_{0}$ be the length of the exploration phase and $q$ be the proportion of buyers in group 0. $\widehat{\bm{\theta}}_{0}$ and $\widehat{\bm{\theta}}_{1}$ are obtained by (2). Under Assumptions 1 and 2, We have with probability at least $1-\eta-(d+2)(e/2)^{-\frac{\lambda_{0}qT_{0}}{2L}}$,
@@
$\|\widehat{\bm{\theta}}_{1}-\bm{\theta}_{1}\|_{2}^{2}\leq\frac{8(d+2)z_{max}^{2}\sigma_{\epsilon}^{2}\log(2/\eta)}{\lambda_{0}^{2}(1-q)T_{0}},$
@@
Recall that $Z_{t}=(z_{1},\cdots,z_{t})^{\top}\in\mathbb{R}^{t\times(d+2)}$ with $z_{t}=(1\ p_{t}\ \bm{x}_{t}^{\top})^{\top}\in\mathbb{R}^{d+2}$ from the exploration phase. We slightly abuse the notation and let $z_{ij}$ be the $(i,j)$-th elements of $Z_{t}$. Under Assumption 1, we have $|z_{ij}|\leq\mathop{\max}\{1,x_{max},B\}:=z_{max}$. Noting that $\epsilon_{i}$ is bounded independent $\sigma_{\epsilon}^{2}$-sub-Gaussian variable with mean zero, and $z_{ij}$ is independent from $\epsilon_{i}$, we know that $z_{ij}\epsilon_{i}$ is zero-mean bounded random variable with variance at most $z_{max}^{2}\sigma_{\epsilon}^{2}$. By Lemma S10, for $0<w_{j}<1$, we have
$~~~~\mathbb{P}\bigg(\bigg|\sum_{i=1}^{t}z_{ij}\epsilon_{i}\bigg|<z_{max}\sigma_{\epsilon}\sqrt{2t\log(2/w_{j})}\bigg)$ (S20)
$=1-\mathbb{P}\bigg(\bigg|\sum_{i=1}^{t}z_{ij}\epsilon_{i}\bigg|\geq z_{max}\sigma_{\epsilon}\sqrt{2t\log(2/w_{j})}\bigg)$
$=1-\mathbb{P}\bigg(\sum_{i=1}^{t}z_{ij}\epsilon_{i}\geq z_{max}\sigma_{\epsilon}\sqrt{2t\log(2/w_{j})}\bigg)-\mathbb{P}\bigg(-\sum_{i=1}^{t}z_{ij}\epsilon_{i}\geq z_{max}\sigma_{\epsilon}\sqrt{2t\log(2/w_{j})}\bigg)$
$\geq 1-w_{j}.$
Let $0<w<1$ and $w_{j}=w/(d+2)$ for $j=1,\cdots,(d+2)$. By (S20), with probability at least 1w1-w, we have
$\bm{\epsilon}_{t}^{\top}Z_{t}Z_{t}^{\top}\bm{\epsilon}_{t}$ $=\sum_{j=1}^{d+2}\bigg(\sum_{i=1}^{t}z_{ij}\epsilon_{i}\bigg)^{2}\leq 2(d+2)z_{max}^{2}\sigma_{\epsilon}^{2}t\log(2/w).$ (S21)
@@
$\|\widehat{\bm{\theta}}-\bm{\theta}\|_{2}^{2}\leq\frac{\bm{\epsilon}_{t}^{\top}Z_{t}Z_{t}^{\top}\bm{\epsilon}_{t}}{\lambda^{2}_{min}(Z_{t}^{\top}Z_{t})}\leq\frac{8(d+2)z_{max}^{2}\sigma_{\epsilon}^{2}\log(2/w)}{\lambda_{0}^{2}t},$ (S22)
with probability at least $1-w-(d+2)(e/2)^{-\frac{\lambda_{0}t}{2L}}$. Since the length of the exploration phase is $T_{0}$, and the proportion of buyers from group 0 is $q$, the numbers of samples used to estimate $\bm{\theta}_{0}$ and $\bm{\theta}_{1}$ are $qT_{0}$ and $(1-q)T_{0}$, respectively. Plugging $t=qT_{0}$ and $t=(1-q)T_{0}$ into (S22), respectively, we can obtain the estimation errors of $\bm{\theta}_{0}$ and $\bm{\theta}_{1}$. ∎
By Lemma S5 with $w=2/T_{0}$, we have with probability at most $\frac{4}{T_{0}}+(d+2)[(e/2)^{-\frac{\lambda_{0}qT_{0}}{2L}}+(e/2)^{-\frac{\lambda_{0}(1-q)T_{0}}{2L}}]$,
@@
$J_{0}=\mathbb{E}(reg_{t})$ $\leq 2B(a_{max}B+b_{max}x_{max})\left\{\frac{4}{T_{0}}+(d+2)[(\frac{e}{2})^{-\frac{\lambda_{0}qT_{0}}{2L}}+(\frac{e}{2})^{-\frac{\lambda_{0}(1-q)T_{0}}{2L}}]\right\}$ (S25)
@@
$p_{1t}=\left\{\begin{aligned} &-\frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}},&\text{if}\ \ \frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}}-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}}\leq\delta-c_{\delta}\sqrt{\frac{\log T_{0}}{T_{0}}},\\ &\widehat{\bm{\gamma}}_{1}^{\top}\widetilde{\bm{x}}_{t}-\delta+\widehat{\gamma}_{2},&\text{if}\ \ \frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}}-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}}>\delta+c_{\delta}\sqrt{\frac{\log T_{0}}{T_{0}}}.\end{aligned}\right$.
@@
$p_{0t}^{\prime}=\left\{\begin{aligned} &-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}},\ &\text{if}\ \ \frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}}-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}}\leq\delta-c_{\delta}\sqrt{\frac{\log T_{0}}{T_{0}}},\\ &\widehat{\bm{\gamma}}_{1}^{\top}\widetilde{\bm{x}}_{t}+\widehat{\gamma}_{2},\ &\text{if}\ \ \frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}}-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}}>\delta+c_{\delta}\sqrt{\frac{\log T_{0}}{T_{0}}}.\end{aligned}\right$.
The regret of our policy depends on the probability $\mathbb{P}(\widehat{\delta}\leq C_{0})$. We define the historical information up to time $t$ as $\mathcal{H}_{t}=\{\bm{x}_{1},\cdots,\bm{x}_{t},G_{1}^{\prime},\cdots,G_{t}^{\prime},y_{1},\cdots,y_{t},p_{1},\cdots,p_{t}\}$. We also define $\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}\cup\{\bm{x}_{t+1},G_{t+1}\}$ as the filtration including $\bm{x}_{t+1}$ and $G_{t+1}$. The expected regret at time $t$ in the exploitation phase is
@@
$=\mathbb{E}\{q\alpha_{0}(p_{0t}^{*2}-p_{0t}^{\prime 2})+(1-q)\alpha_{1}(p_{1t}^{*2}-p_{1t}^{2})+q\bm{\beta}_{0}^{\top}\widetilde{\bm{x}}_{t}(p_{0t}^{*}-p_{0t}^{\prime})+(1-q)\bm{\beta}_{1}^{\top}\widetilde{\bm{x}}_{t}(p_{1t}^{*}-p_{1t})\}$
@@
$\leq a_{max}\mathbb{E}(p_{0t}^{*}-p_{0t}^{\prime})^{2},$
where the fourth equality is from $p_{1t}^{*}=p_{0t}^{*}-\delta$ and $p_{1t}=p_{0t}^{\prime}-\delta$, and the last second equality is due to $2(q\alpha_{0}+(1-q)\alpha_{1})p_{0t}^{*}=2(1-q)\alpha_{1}\delta-(q\bm{\beta}_{0}+(1-q)\bm{\beta}_{1})^{\top}\widetilde{\bm{x}}_{t}$ from (3). We now upper bound the price difference between the optimal policy and our policy. By (4), we rewrite the pricing parameters as
@@
[…]
@@
$\leq T_{0}\bar{B}+\frac{c^{\prime}_{4}T}{T_{0}}+2\sum_{t=T_{0}}^{T}(\frac{c^{\prime}_{4}}{T_{0}}+q\bar{B})\eta_{t-1}$
@@
$=\sqrt{\left[\frac{c^{\prime\prime}_{2}(d+2)+c^{\prime\prime}_{3}}{q(1-q)}\right]T}+(c^{\prime\prime}_{4}+c^{\prime\prime}_{5}q)\sum_{t=2}^{T}\eta_{t-1}$
for some positive constants $c^{\prime\prime}_{2},c^{\prime\prime}_{3},c^{\prime\prime}_{4}$ and $c^{\prime\prime}_{5}$. To minimize the number of constants, we have
$Regret_{T}\leq\sqrt{\frac{c_{2}dT}{q(1-q)}}+c_{3}q\sum_{t=2}^{T}\eta_{t-1}$
for some positive constants $c_{2}=\frac{c^{\prime\prime}_{2}(d+2)+c^{\prime\prime}_{3}}{d}$ and $c_{3}=\frac{c^{\prime\prime}_{4}+c^{\prime\prime}_{5}q}{q}$.
S.9.6 Proof of Theorem 3
@@
    The third equality is from $p_{0}^{*}-p_{1}^{*}=\delta$ and $p_{0}-p_{1}=\delta$. The fourth equality is due to $\delta=1/4$ and $6\alpha p_{0}^{*}=\frac{7\alpha}{2}-1$ derived from $p_{0}^{*}=\frac{7}{12}-\frac{1}{6\alpha}$. the last line is from $\alpha\in[-1/2,-1/5]$. Similarly, we can obtain $R_{0}(p_{0}^{*},\alpha)-R_{0}(p_{0},\alpha)+R_{1}(p_{1}^{*},\alpha)-R_{1}(p_{1},\alpha)\geq\frac{3}{5}(p_{1}^{*}-p_{1})^{2}$.
@@
$\mathbf{X}_{k}\succeq\bm{0}\ \textit{and}\ \lambda_{max}(\mathbf{X}_{k})\leq L\ \textit{almost surely.}$